OpenStudy (anonymous):

square both sides of the equation: x = 5 + 3y

OpenStudy (anonymous):

\[(x)^2=(5+3y)^2\] \[(x+y)^2=x^2+2xy+y^2\] \[\therefore x^2=(5)^2+2(5)(3y)+(3y)^2\] \[x^2=25+30y+9y^2\]